# NON-CLASSICAL GENERATING SETS FOR FUCHSIAN SCHOTTKY GROUPS

### v6

@@ 3. PROOF OF OUR MAIN THEOREM
Recall that, we have intended to prove this theorem by contradiction. Let, $SC_{\kappa^{*}}$ be a semi-circle meeting the closed interval $[-(\lambda+2)(\lambda+3),$ $-(\lambda+3)]$ which is analogous to the semi-circle $SC_{4,\kappa}$ for the group generated by $\Big{(}h_{(4,1)_{\kappa}}^{**}\Big{)}^{2}$. The semi-circle $SC_{4,\kappa}$ separates the fixed points $\tau$ and $\{-\tau\}$ of the transformation ${h^{**}}_{(4,1)_{\kappa}}$. So, the semi-circle $SC_{\kappa^{*}}$ also separates these fixed points. Since the semi-circle $SC_{\kappa^{*}}$ is outside the semi-circle $SC_{3,\kappa}$, its image ${h^{*}}_{(3,2)_{\kappa}}(SC_{\kappa^{*}})$ meets the closed interval $[(\lambda-1),$ $(\lambda+1)]$ twice. Again, the fixed point $\tau$ and the point $\Big{\{}-(\lambda+2)(\lambda+3)\Big{\}}$ are outside of the semi-circle $SC_{\kappa^{*}}$. Hence, the points ${h^{*}}_{(3,2)_{\kappa}}(\tau)$ and ${h^{*}}_{(3,2)_{\kappa}}\Big{\{}-(\lambda+2)(\lambda+3)\Big{\}}$ will be separated by the semi-circular curve ${h^{*}}_{(3,2)_{\kappa}}(SC_{\kappa^{*}})$ and placed within this curve by situating on the boundary of the hyperbolic plane. Therefore, the diameter of ${h^{*}}_{(3,2)_{\kappa}}(SC_{\kappa^{*}})$ will be greater than $\frac{1}{5}$, since
@@
But Lemma 2.5 shows that the gaps between the curves in the set ${SC}^{*}_{\kappa}$ are less than $\frac{1}{5}$. So, ${h^{*}}_{(3,2)_{\kappa}}(SC_{\kappa^{*}})$ intersects some of the semi-circular curves $SC_{\kappa}^{j}$ (where $j=1,2,3,...,\textbf{I}^{+}$) in the family ${SC}^{*}_{\kappa}$. Hence, all images of the semi-circles $SC_{1,\kappa},SC_{2,\kappa},$ $SC_{3,\kappa}$, and $SC_{4,\kappa}$ are not disjoint, which is impossible. So we reach a contradiction. Therefore, the Schottky group $\Gamma^{N}_{S_{\kappa}}$ $=$ $<{h^{*}}_{(3,2)_{\kappa}},{h^{**}}_{(4,1)_{\kappa}}>$ is a non-classical Fuchsian Schottky group. This completes the proof of Theorem 1. ∎
@@ 4. EXAMPLES OF FUCHSIAN SCHOTTKY GROUPS WITH NON-CLASSICAL GENERATING SETS
[…]

### v7

@@ 4. PROOF OF OUR MAIN THEOREM
Recall that, we have intended to prove this theorem by contradiction. Let, $SC_{\kappa^{*}}$ be a semi-circle meeting the closed interval $(\lambda+2)(\lambda+3),$ $(\lambda+3)]$ which is analogous to the semi-circle $SC_{4,\kappa}$ for the group generated by $\Big(h_{(4,1)_{\kappa}}^{**}\Big)^{2}$. The semi-circle $SC_{4,\kappa}$ separates the fixed points $\tau$ and $\{-\tau\}$ of the transformation ${h^{**}}_{(4,1)_{\kappa}}$. So, the semi-circle $SC_{\kappa^{*}}$ also separates these fixed points. Since the semi-circle $SC_{\kappa^{*}}$ is outside the semi-circle $SC_{3,\kappa}$, its image ${h^{*}}_{(3,2)_{\kappa}}(SC_{\kappa^{*}})$ meets the closed interval $(\lambda-1),$ $(\lambda+1)]$ twice. Again, the fixed point $\tau$ and the point $\Big\{-(\lambda+2)(\lambda+3)\Big\}$ are outside of the semi-circle $SC_{\kappa^{*}}$. Hence, the points ${h^{*}}_{(3,2)_{\kappa}}(\tau)$ and ${h^{*}}_{(3,2)_{\kappa}}\Big\{-(\lambda+2)(\lambda+3)\Big\}$ will be separated by the semi-circular curve ${h^{*}}_{(3,2)_{\kappa}}(SC_{\kappa^{*}})$ and placed within this curve by situating on the boundary of the hyperbolic plane. Therefore, the diameter of ${h^{*}}_{(3,2)_{\kappa}}(SC_{\kappa^{*}})$ will be greater than $\frac{1}{5}$, since
@@
But Lemma 3.5 shows that the gaps between the curves in the set ${SC}^{*}_{\kappa}$ are less than $\frac{1}{5}$. So, ${h^{*}}_{(3,2)_{\kappa}}(SC_{\kappa^{*}})$ intersects some of the semi-circular curves $SC_{\kappa}^{j}$ (where $j=1,2,3,...,\textbf{I}^{+}$) in the family ${SC}^{*}_{\kappa}$. Hence, all images of the semi-circles $SC_{1,\kappa},SC_{2,\kappa},$ $SC_{3,\kappa}$, and $SC_{4,\kappa}$ are not disjoint, which is impossible. So we reach a contradiction. Therefore, the Schottky group $\Gamma^{N}_{S_{\kappa}}$ == ${h^{*}}_{(3,2)_{\kappa}},{h^{**}}_{(4,1)_{\kappa}}>$ is a non-classical Fuchsian Schottky group. This completes the proof of Theorem 1. ∎
@@ 5. EXAMPLES OF FUCHSIAN SCHOTTKY GROUPS WITH NON-CLASSICAL GENERATING SETS
$\text{ and }{h^{4}}_{(4,1)_{\kappa_{2}}}:z$ $\longrightarrow$ $\frac{(2+\mu)(1-\kappa_{2})^{-1}z+(1-\kappa_{2})\{(2+\mu)^{2}(1-\kappa_{2})^{-2}-1\}}{(1-\kappa_{2})^{-1}+(2+\mu)(1-\kappa_{2})^{-1}}$ (17)